# Low-Frequency Black-Box Backdoor Attack via Evolutionary Algorithm

### 23 Feb 2024 v1

@@ Abstract
While convolutional neural networks (CNNs) have achieved success in computer vision tasks, it is vulnerable to backdoor attacks. Such attacks could mislead the victim model to make attacker-chosen prediction with a specific trigger pattern. Until now, the trigger injection of existing attacks is mainly limited to spatial domain. Recent works take advantage of perceptual properties of planting specific patterns in the frequency domain, which only reflect indistinguishable pixel-wise perturbations in pixel domain. However, in the black-box setup, the inaccessibility of training process often renders more complex trigger designs. Existing frequency attacks simply handcraft the magnitude of spectrum, introducing anomaly frequency disparities between clean and poisoned data and taking risks of being removed by image processing operations (such as lossy compression and filtering). In this paper, we propose a robust low-frequency black-box backdoor attack (LFBA), which minimally perturbs low-frequency components of frequency spectrum and maintains the perceptual similarity in spatial space simultaneously. The key insight of our attack restrict the search for the optimal trigger to low-frequency region that can achieve high attack effectiveness, robustness against image transformation defenses and stealthiness in dual space. We utilize simulated annealing (SA), a form of evolutionary algorithm, to optimize the properties of frequency trigger including the number of manipulated frequency bands and the perturbation of each frequency component, without relying on the knowledge from the victim classifier. Extensive experiments on real-world datasets verify the effectiveness and robustness of LFBA against image processing operations and the state-of-the-art backdoor defenses, as well as its inherent stealthiness in both spatial and frequency space, making it resilient against frequency inspection.
@@ 1 Introduction
CNNs are vulnerable to backdoor attacks [1, 5, 15, 32, 33, 35]. Such attacks can mislead the model to make attack-chosen predictions with triggers in the use phase while behaving normally on clean images, which causes severe consequences in high-stakes applications such as autonomous driving [14] and biometric authentication [19]. For example, in a face recognition model with a backdoor, the presence of a specific trigger always leads to the identification of a face as Donald Trump.
@@
Prior backdoor attacks possess the capability to inject imperceptible triggers into spatial domain [10, 30, 36, 9, 55, 54, 6]. However, inserting triggers in spatial space always harm the semantics of infected image pixels (see Figure 1). Recent works have revealed that backdoor attacks can inject trigger patterns into frequency space [49, 12, 17, 20]. For example, FTrojan [49] manipulates the frequency spectrum of input images with a specfic perturbation within fixed frequency band. However, manually crafting frequency components, especially in high-frequency regions, could harm the robustness of trigger and thus trigger effectiveness can be removed by image processing operations such as lowpass filters. Additionally, both spatial and frequency triggers introduce distinguishable artifacts when transformed to frequency space (see Figure 1 and Figure 5).
@@
Extending from [23], an ideal black-box backdoor attack should achieve four objectives, namely, functionality-preserving, effectiveness, dual-space stealthiness, and robustness. Functionality-preserving ensures high test accuracy on clean data. Effectiveness is demonstrated by the ability to misclassify poisoned data to the target label with a high probability. Dual-space stealthiness implies that poisoned images exhibits visual and frequency similarity to clean ones. Robustness is demonstrated by its effectiveness against image transformations and resistance to backdoor defenses. Although [23] successfully achieved the goals at pixel level, they do not consider the stealthiness in the frequency perspective. An efficient solution that satisfies all requirements is to design stealthy triggers in the frequency domain. Typically, low-frequency components of an image contains essential information, while the high-frequency components capture finer details and noise. According to prior work [7], inserting triggers in low-frequency region offers the following advantages: (1) low-frequency components have a perceptual capacity that allows trigger insertion without perceptual degradation; (2) low-frequency components exhibit greater resilience in lossy compression operations such as JPEG, whereas high-frequency components are more pronounced to data loss; and (3) trigger inserted in low-frequency region is harder to be removed by low-pass filtering compared to the high-frequency region.
@@
Building on the insights mentioned above, we introduce LFBA, a robust low-frequency backdoor attack with imperceptible triggers in dual space. The key insight of our LFBA is to find the optimal trigger that can achieve both dual-space stealthiness and attack robustness when the model and defense strategy are unknown. Our work marks the first exploration into the robustness from a frequency perspective, which searches mininal perturbations in low-frequency region against image transformations while naturally guaranteeing perceptual similarity in the pixel domain. However, finding an appropriate frequency trigger is non-trivial: a large perturbation can disrupt invisibility of poisoned images, while a small perturbation may hinder the model’s ability to learn the trigger feature, resulting in low attack effectiveness. Additionally, selecting an improper frequency band for trigger insertion can compromise the attack robustness against image processing operations (see Table 3). For example, malicious perturbations in high-frequency regions can be mitigated by lowpass filters. To address these challenges in a black-box scenario, we leverage simulated annealing (SA), a gradient-free optimziation algorithm, to effectively search for the optimal trigger in frequency domain. Specifically, we first convert a clean image to frequency domain with discrete cosine transform (DCT). Then, we iteratively optimize the perturbations and frequency bands of trigger with SA, in order to maximize the attack effectiveness as the primary goal and a penalty term concerning the dual-space stealthiness. Due to lacking knowledge about the target model in the black-box setting, we approximate the quality of trigger in terms of backdoor loss in the optimization process with a semi-trained evaluation model. Finally, we use inverse DCT (IDCT) to produce the poisoned image. Since the imperceptible perturbations is posed in frequency space, our proposed LFBA preserves the invisibility in the spatial domain and is robust to any eradication in frequency domain. The main contributions in this work are summarized as follows:
@@
$\bullet$ We explore both stealthiness and robustness of trigger from frequency perspective and propose an imperceptible and robust black-box backdoor attack in dual space.
@@
$\bullet$ We propose a constrained optimization problem to find optimal frequency triggers. Then we utilize simulated annealing, an efficient evolutionary algorithm, to solve the objective without relying on gradient information.
@@ 2 Related Work
Backdoor Attacks. Work [15] introduced the first backdoor attack against deep learning models. They employ a patch-based pattern as trigger, injecting it into a small fraction of clean data, which causes the victim model to misclassify those poisoned images to the target label. After that, various attacks, aiming at improving stealthiness and robustness through the design of triggers and training process, have been proposed in the literature. Current backdoor attacks can be roughly categorized into two main aspects: (1) spatial domain-based attacks and (2) frequency domain-based attacks.
@@
(1) Spatial domain-based attacks: To enhance the invisibility of triggers for bypassing human inspection, some works [1, 33, 10, 30, 36, 23] focus on the imperceptible backdoor attacks in spatial domain. For example, [1] uses sinusoidal signals as triggers which results in slightly varying backgrounds. [33] utilizes the natural reflection as trigger into the victim model. [30] leverages DNN-based image steganography technique to hide an attacker-specified string into clean images as sample-specific triggers. Later, several works [9, 55, 54, 6] reveal the importance of stealthiness in latent feature space. For example, [9] learns a trigger generator to constrain the similarity of hidden features between poisoned and clean data via Wasserstein regularization. Further, to improve the stealthiness of triggers, [54] adaptively learns the generator by constraining the latent layers, which makes triggers more invisible in both input and latent feature space. Although the above spatial attacks provide promising natural stealthiness, they did not consider the robustness against image processing operations which are commonly used in data preprocessing period and thus their attack effectiveness will be heavily harmed by such operations. Moreover, most attacks need a strong attack assumption that the adversary should have full control of training process and the knowledge of the victim model. More importantly, many spatial backdoor attacks exhibit severe high-frequency artifacts that could be easily detected in the frequency domain (see Figure 1 and Figure 5).
@@
(2) Frequency domain-based attacks: Recent works [17, 49, 12, 51, 20] explore another attack surface, namely, frequency domain, which naturally guarantee invisibility due to the properties of frequency components. For example, [49] handcrafts two single frequency bands with fixed perturbations as trigger. [12] injects the low-frequency information of a trigger image by linearly combining the spectral amplitude of poisoned and clean images. However, they all introduce distinguishable frequency artifacts and can be detected via frequency inspection. Moreover, they also focus on natural stealthiness and do not consider robustness against image processing operations. Different from the above works, we propose the first black-box frequency backdoor attack that achieves imperceptibility in dual space and robust against image processing defenses.
@@
Backdoor Defenses. Existing defenses can be roughly divided into two categories, i.e., defensive [31, 48, 4, 38, 53, 28, 29] and detective [13, 3, 45, 24, 51] mechanisms. Defensive methods focus on mitigating the effectiveness of potential backdoor attacks. For example, fine-pruning [31] prunes the dormant neurons in the last convolution layer based on clean inputs’ activation values. Neural Cleanse [48] reconstructs potential triggers for each target label via reverse engineering and renders the backdoor ineffective by retrain patches strategy. Neural Attention Distillation [28] uses a “teacher" model to guide the finetuning of the backdoored “student" network to erase backdoor triggers. Representative detective methods include STRIP [13] that perturbs or superimposes clean inputs to identify the potential backdoors during inference time, spectral signature [45] that utilizes latent feature representations to detect outliers and [51] that uses supervised learning to differentiate between clean and poisoned data in frequency space. Besides, image processing-based methods [29, 49, 39] have also been studied , which remove backdoors by image processing transformations.
In this work, we showcase that our proposed attack can evade representative defenses including frequency inspection, image processing operations and mainstream backdoor defenses.
@@
Threat Model. We consider the same black-box scenario as in prior works [40, 23, 49] where the adversary, i.e. a malicious data provider, can only inject a limited number of poisoned samples into clean training set for public use. However, the attacker should not have control of the training process or the knowledge of the victim model, which is a more practical and challenging attack scenario than white-box attacks [6, 9, 10, 11, 54]. Hence, the only task of our attack is to design the trigger.
@@ 3.1 Preliminaries
We consider backdoor attacks on image classification. Let $f_{\theta}:\mathcal{I}\rightarrow\mathbb{R}^{K}$ be an image classifier parameterized with $\theta$ that maps an input image $\mathcal{I}\subseteq[0,1]^{H\times W\times C}$ to a output class, where $K$ is the number of classes, $H$, $W$ and $C$ are the height, width and channels of an input image. The parameters $\theta$ of the classifier are learned using a training dataset $D_{c}=\{(x_{i},y_{i})|x_{i}\in\mathcal{I},y_{i}\in\mathbb{R}^{K}\}^{N}_{i=1}$.
@@ 3.2 Frequency Backdoor Attack
We redesign the trigger function $\mathcal{T}$ in the frequency space to better search the frequency trigger that can achieve dual-space stealthiness. Given a clean sample $(x,y)$ in $D_{c}$, we first transform it to the frequency domain via DCT function $\mathcal{D(\cdot)}$ and obtain the frequency form $x^{f}$ of input $x$ as:
$x^{f}(h^{f},w^{f},c)$ $=\mathcal{D}(x(h,w,c))$ (3)
$=w(h^{f})w(w^{f})\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\sum_{c=0}^{C-1}x(h,w,c)\cos{\left[\frac{(2h+1)h^{f}\pi}{2H}\right]}\cos{\left[\frac{(2w+1)w^{f}\pi}{2W}\right]}$
for $\forall h^{f}=0,1,...,H-1$ and $\forall w^{f}=0,1,...,W-1$, where $H,W,C$ represent the height, width and number of channels of the image. For simplicity, we assume $H=W$, therefore $w(0)=\sqrt{\frac{1}{4H}}$ and $w(k)=\sqrt{\frac{1}{2H}}$ for $k>0$. Accordingly, $\mathcal{D}^{-1}(\cdot)$ denotes the IDCT. Therefore, our frequency trigger function is defined as follows:
@@
Our primary goal is to search the optimal trigger that can achieve high attack effectiveness. Ideally, the effectiveness of trigger should be assessed by using the victim classifier, whereas the adversary lacks knowledge about it in practical attack scenario. As a alleviation, we utilize the poisoned set to learn a semi-trained evaluation classifier $f^{s}$, which may be derived from open-source architectures. This evaluation classifier can approximate the trigger effectiveness with an acceptable deviation in practice. Therefore, the main task of our attack is defined as follows:
@@
One may argue that large crafted perturbations in specific frequency bands could also achieve high attack effectiveness and practical natural stealthiness such as [49, 12]. However, frequency triggers without careful consideration may cause inconsistency in the frequency space (see Figure 1 and Figure 5). We hereby use dual-space stealthiness penalties to ensure imperceptible perturbations and invisibility of poisoned images in both frequency and spatial space. We define penalty term for dual-space stealthiness as follows [^1] The distance between clean and poisoned samples remains consistent between frequency and spatial domains. We hereby only constrain frequency disparities.:
@@
where $p=2$ denoting $l_{2}$-norm distance used to calculate the disparity between clean and poisoned sample in the frequency domain. Taking into account both the primary goal and the dual-space stealthiness penalty, our attack goal aims to minimize the overall objective function under the constraint w.r.t. frequency perturbation. The optimization problem is formulated as follows:

### 06 Mar 2024 v2

@@ Abstract
Convolutional Neural Networks (CNNs) that have excelled in diverse computer vision tasks are vulnerable to backdoor attacks, enabling attacker-controlled predictions via specific patterns. Restricted to spatial domains, recent research exploits perceptual traits by embedding patterns in the frequency domain, yielding pixel-level indistinguishable perturbations. In black-box settings, restricted access to training and model necessitates advanced trigger designs. Current frequency-based attacks manipulate magnitude spectra, introducing discrepancies between clean and poisoned data, though vulnerable to common image processing operations like compression and filtering.
In this paper, we propose a robust low-frequency practical backdoor attack (LFBA) in black-box setup that minimally perturbs low-frequency components of frequency spectrum and maintains the perceptual similarity in spatial space simultaneously. Our methodology capitalizes on the insight that optimal triggers can be located in low-frequency regions to maximize attack effectiveness, robustness against image transformation defenses, and stealthiness in dual space. We utilize simulated annealing (SA), a form of evolutionary algorithm, to optimize the properties of frequency trigger including the number of manipulated frequency bands and the perturbation of each frequency component, without relying on prior knowledge of the victim classifier. Extensive experiments on real-world datasets confirm the effectiveness and robustness of LFBA against image processing operations and state-of-the-art backdoor defenses. Furthermore, LFBA exhibits inherent stealthiness in both spatial and frequency spaces, making it resistant to frequency inspection.
@@ 1 Introduction
CNNs are vulnerable to backdoor attacks [1, 5, 15, 32, 33, 35] that can mislead the model to make attack-chosen predictions with triggers in the use phase while behaving normally on clean images, causing severe consequences in high-stakes applications such as autonomous driving [14] and biometric authentication [19].
@@
Prior backdoor attacks possess the capability to inject imperceptible triggers into spatial domain [10, 30, 36, 9, 55, 54, 6]. Inserting triggers in spatial space can harm the semantics of infected image pixels (see Figure 1). Recent works have concluded that backdoor attacks can inject trigger patterns into frequency space [49, 12, 17, 20]. For example, FTrojan [49] manipulates mid- and high-frequency spectrum of input images with a pre-defined perturbation within fixed frequency band. Manually crafting frequency components, especially in high-frequency regions, could harm the robustness of trigger and thus trigger effectiveness can be eliminated by image processing operations such as lowpass filters. Moreover, both spatial and frequency triggers introduce distinguishable artifacts when transformed to frequency space (see Figure 1 and Figure 5).
@@
Inspired by [23], an ideal and practical backdoor attack should achieve four objectives, namely, functionality preservation, effectiveness, dual-space stealthiness, and robustness. Functionality preservation ensures high test accuracy on clean data. Effectiveness is demonstrated by the ability to misclassify poisoned data to the target label with a high probability. Dual-space stealthiness implies that poisoned images exhibit visual and frequency similarity to clean ones. Robustness is demonstrated by its effectiveness against image transformations and resistance to backdoor defenses. Although successfully achieving the goals at pixel level, [23] does not consider the stealthiness in the frequency perspective. This work explores a new perspective of attacks in the frequency domain.
Typically, low-frequency components of an image contain semantic information, while the high-frequency components capture finer details and noise. According to prior works such as [7, 16], inserting triggers in low-frequency region offers concrete influences: (1) low-frequency components have a perceptual capacity that allows trigger insertion without perceptual degradation; (2) low-frequency components exhibit greater resilience in lossy compression operations such as JPEG, whereas high-frequency components are more pronounced to data loss; and (3) trigger inserted in low-frequency region is harder to be removed by low-pass filtering compared to the high-frequency region.
@@
Building on the above insights, we develop LFBA, a robust and practical low-frequency backdoor attack with imperceptible triggers in dual space. The key insight of LFBA is to find the optimal trigger that can achieve both dual-space stealthiness and attack robustness when the model and defense strategy are unknown. This design marks the first exploration into the robustness from a frequency perspective, which locates mininal perturbations in low-frequency region against image transformations while naturally guaranteeing perceptual similarity in the pixel domain. Finding such a frequency trigger in a black-box scenario is not trivial. Due to the absence of the victim model and training process, one may handcraft the frequency trigger, but it could lead to improper frequency properties of the trigger. For example, a large perturbation can disrupt invisibility of poisoned images, while a small perturbation could hinder the model’s ability to learn the trigger feature, resulting in low attack effectiveness. Furthermore, selecting an improper frequency band for trigger insertion can compromise the attack robustness against image processing operations (see Table 3).
To address the challenges, we leverage simulated annealing (SA), a gradient-free optimziation algorithm, to effectively search for the optimal trigger in frequency domain. Specifically, we first convert a clean image to frequency domain with discrete cosine transform (DCT). We iteratively optimize the perturbations and frequency bands of trigger with SA, in order to maximize the attack effectiveness as the primary goal and a penalty term concerning the dual-space stealthiness. Due to lacking knowledge about the target model in the black-box setting, we approximate the quality of trigger in terms of backdoor loss in the optimization process with a semi-trained evaluation model. Finally, we use inverse DCT (IDCT) to produce the poisoned image. Since the imperceptible perturbations is posed in frequency space, LFBA preserves the invisibility in the spatial domain and is robust to any eradication in frequency domain. The main contributions of this work are as follows:
@@
$\bullet$ We explore both stealthiness and robustness of trigger from frequency perspective and design an imperceptible and robust black-box backdoor attack in dual space.
@@
$\bullet$ We propose a constrained optimization problem to find optimal frequency triggers. Then we utilize SA to solve the objective without relying on gradient information.
@@ 2 Related Work
Backdoor Attacks. [15] introduces the first backdoor attack against deep learning models. It employs a patch-based pattern as trigger, injecting it into a small fraction of clean data, which causes the victim model to misclassify those poisoned images to the target label. After that, various attacks, aiming at improving stealthiness and robustness through the design of triggers and training process, have been proposed in the literature.
@@
(1) Spatial domain-based attacks: To enhance the invisibility of triggers for bypassing human inspection, some works [1, 33, 10, 30, 36, 23] focus on imperceptible backdoor attacks in spatial domain. For example, [1] uses sinusoidal signals as triggers which results in slightly varying backgrounds; [33] utilizes natural reflection as trigger into the victim model, while [30] leverages DNN-based image steganography technique to hide an attacker-specified string into clean images as sample-specific triggers. Later, several works [9, 55, 54, 6] reveal the importance of stealthiness in latent feature space. [9] learns a trigger generator to constrain the similarity of hidden features between poisoned and clean data via Wasserstein regularization. To improve the stealthiness of triggers, [54] adaptively learns the generator by constraining the latent layers, which makes triggers more invisible in both input and latent feature space. While spatial attacks offer inherent stealthiness, they often overlook robustness against common image processing operations utilized during data preprocessing. Consequently, their effectiveness is significantly compromised by such operations. Moreover, most attacks require a strong attack assumption that the adversary possesses full control over the training process and has knowledge of the victim model. More importantly, many spatial backdoor attacks exhibit severe high-frequency artifacts that can be easily detected in the frequency domain (see Figure 1 and Figure 5).
@@
(2) Frequency domain-based attacks: Recent works [17, 49, 12, 51, 20] explore another attack surface, namely, frequency domain, naturally guaranteeing invisibility due to frequency properties. [49] handcrafts two single frequency bands with fixed perturbations as trigger, and [12] injects low-frequency information of a trigger image by linearly combining spectral amplitude of poisoned and clean images. These works introduce distinguishable frequency artifacts and can be detected via frequency inspection. Moreover, they focus on natural stealthiness and do not consider robustness against image transformation operations.
Different from the above works, we propose a black-box frequency backdoor attack that firstly achieves imperceptibility in dual spaces and robust against image processing defenses.
@@
Backdoor Defenses. Defensive [31, 48, 4, 38, 53, 28, 29] and detective [13, 3, 45, 24, 51] mechanisms are commonly used for backdoor defenses. Defensive methods focus on mitigating the effectiveness of potential backdoor attacks. For example, fine-pruning [31] prunes the dormant neurons in the last convolution layer based on clean inputs’ activation values. Neural Cleanse [48] reconstructs potential triggers for each target label via reverse engineering and renders the backdoor ineffective by retrain patches strategy. Neural Attention Distillation [28] uses a “teacher" model to guide the finetuning of the backdoored “student" network to erase backdoor triggers. Representative detective methods include STRIP [13] which perturbs or superimposes clean inputs to identify the potential backdoors during inference time, spectral signature [45] using latent feature representations to detect outliers and [51] leveraging supervised learning to differentiate between clean and poisoned data in frequency space. Besides, image processing-based methods [29, 49, 39] have been studied, which remove backdoors by image processing transformations. In this work, we showcase that the proposed attack can evade representative defenses including frequency inspection, image processing operations and mainstream backdoor defenses.
@@
Threat Model. We consider rather realistic black-box scenario as in prior works [40, 23, 49] where the adversary, i.e. a malicious data provider, can only inject a limited number of poisoned samples into clean training set for public use. The attacker should not have control over the training process or have knowledge of the victim model. This is a more practical and challenging attack scenario than white-box attacks [6, 9, 10, 11, 54].
@@ 3.1 Preliminaries
We consider backdoor attacks on image classification. Let $f_{\theta}:\mathcal{I}\rightarrow\mathbb{R}^{K}$ be an image classifier parameterized with $\theta$ that maps an input image $\mathcal{I}\subseteq[0,1]^{H\times W\times C}$ to an output class, where $K$ is the number of classes, $H$, $W$ and $C$ are the height, width and channels of an input image. The parameters $\theta$ of the classifier are learned using a training dataset $D_{c}=\{(x_{i},y_{i})|x_{i}\in\mathcal{I},y_{i}\in\mathbb{R}^{K}\}^{N}_{i=1}$.
@@ 3.2 Frequency Backdoor Attack
We redesign the trigger function $\mathcal{T}$ in the frequency space to better search the frequency trigger that can achieve dual-space stealthiness. Given a clean sample $(x,y)$ in $D_{c}$, we first transform it to the frequency domain via DCT function $\mathcal{D(\cdot)}$ and obtain the frequency form $x^{f}(h^{f},w^{f},c)$ of input $x(h,w,c)$ as:
$\mathcal{D}(x(h,w,c))=V(h^{f})V(w^{f})\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}\sum_{c=0}^{C-1}x(h,w,c)\cos{\left[\frac{(2h+1)h^{f}\pi}{2H}\right]}\cos{\left[\frac{(2w+1)w^{f}\pi}{2W}\right]}$ (3)
for $\forall h^{f}=0,1,...,H-1$ and $\forall w^{f}=0,1,...,W-1$, where $H,W,C$ represent the height, width and number of channels of the image. For simplicity, we assume $H=W$, therefore $V(0)=\sqrt{\frac{1}{4H}}$ and $V(k)=\sqrt{\frac{1}{2H}}$ for $k>0$. Accordingly, $\mathcal{D}^{-1}(\cdot)$ denotes the IDCT as follows:
$\mathcal{D}^{-1}(x^{f}(h^{f},w^{f},c))=\sum_{h^{f}=0}^{H-1}\sum_{w^{f}=0}^{W-1}\sum_{c=0}^{C-1}V(h)V(w)x^{f}(h^{f},w^{f},c)\cos{\left[\frac{(2h^{f}+1)h\pi}{2H}\right]}\cos{\left[\frac{(2w^{f}+1)w\pi}{2W}\right]}$ (4)
Therefore, our frequency trigger function is defined as:
@@
Our primary goal is to search the optimal trigger that can achieve high attack effectiveness. Ideally, the effectiveness of trigger should be assessed by using the victim classifier, whereas the adversary lacks knowledge about it in practical attack scenario. As an alleviation, we utilize the poisoned set to learn a semi-trained evaluation classifier $f^{s}$, which may be derived from open-source architectures. This evaluation classifier can approximate the trigger effectiveness with an acceptable deviation in practice. Therefore, the main task of our attack to minimize is defined as follows:
@@
One may argue that large crafted perturbations in specific frequency bands could also achieve high attack effectiveness and practical natural stealthiness such as [49, 12]. Frequency triggers without careful consideration can bring distinguishable artifacts in the frequency space (see Figure 1 and Figure 5). We hereby use dual-space stealthiness penalties to ensure imperceptible perturbations and invisibility of poisoned images in both frequency and spatial space. We define penalty term for dual-space stealthiness as follows [^1] The distance between clean and poisoned samples remains consistent between frequency and spatial domains. We hereby only constrain frequency disparities.:
@@
where $p=2$ denotes $l_{2}$-norm distance used to calculate the disparity between clean and poisoned sample in the frequency domain. Taking into account both the primary goal and the dual-space stealthiness penalty, our attack goal aims to minimize the overall objective function under the constraint w.r.t. frequency perturbation. The optimization problem is formulated as follows: